An equilateral triangle with side length 8cm has its base on the x-axis and top vertex on the y-axis. A rectangle is inscribed in the triangle. Determine the largest area of the rectangle.
1
Expert's answer
2020-06-10T19:56:20-0400


The area of the rectangle:

"S=2xy"

We have:

"y=h-xtan60\degree"

where h is height of triangle

"y=a\sqrt{3}/2-x\sqrt{3}"

where a is side length

Then:

"S=2x(a\sqrt{3}/2-x\sqrt{3})"

"\frac {dS}{dx}=a\sqrt{3}-4x\sqrt{3}=0"

"x=a\/4"

"S_{max}=2\frac{a}{4}(\frac{a\sqrt{3}}{2}-\frac{a\sqrt{3}}{4})=\frac{a^2\sqrt{3}}{8}"

Answer:

"S_{max}=\frac{8^2\sqrt{3}}{8}=8\sqrt{3}" cm